OpenStudy (valerie):

Joe Jay purchased a new colonial home for $260,000, putting down 20%. He decided to use Loyal Bank for his mortgage. They were offering a 6 1/2% for a 25-year mortgage. The principal after the first payment had a balance outstanding of:

OpenStudy (valerie):

A. $207,270.95 B. $207,720.59 C. $207,720.95 D. $207,270.59

OpenStudy (wolf1728):

The loan is for 260,000 minus 20% of 260,000 = 52,000 So, loan is for 208,000

OpenStudy (valerie):

How do i calculate the principal after the first payment?

OpenStudy (wolf1728):

Go here to see the loan payment formula: http://www.1728.org/loanform.htm And here is how the interest, equity and balance of a loan are calculated: http://www.1728.org/loanmath.htm

OpenStudy (valerie):

I keep getting 206,595.57 but that's not one of the options?

OpenStudy (wolf1728):

First of all do you know the monthly payment?

OpenStudy (valerie):

1404.43

OpenStudy (wolf1728):

For the monthly payment I calculate 1,372.10 *************************************************** Well we have to convert the annual rate to a monthly rate 6.25/1200 = 0.0052083333 First payment interest = 208,000 * 0.0052083333 1,083.33 Equity = Loan Pmt - Interest Equity = 1,372.10 -1,083.33 288.77 Loan balance = Previous Balance - Equity Loan balance = 208,000 -288.77 207,711.23 And that is NOT one of the choices either

OpenStudy (valerie):

okay it's a weird question. I think i'll guess

OpenStudy (wolf1728):

Okay. Those 2 links I gave you should provide a little help. Okay and good luck :-)